Use the discriminant to determine how many real-number solutions the equation has.
1 - 7a2 = -7a - 2

Answer by algebrahouse.com(1079)

"Use the discriminant to determine how many real-number solutions the equation has.
1 - 7a2 = -7a - 2"

Discriminant is b² - 4ac
If b² - 4ac > 0, there are two real number solutions
If b² - 4ac = 0, there is one real number solution
If b² - 4ac < 0, there are two imaginary number solutions

1 - 7a² = -7a - 2
7a² - 7a - 3 = 0 {added 7a² and subtracted 1 from both sides}
a = 7, b = -7, c = -3

b² - 4ac
= (-7)² - 4(7)(-3)
= 49 + 84
= 133

133 > 0,there are two real number solutions
www.algebrahouse.com